OpenStudy (anonymous):

Solve the following system by substitution y=x-1 y=-4/3x-8 A. 4,3 B. 3,4 C. -4,-3 D.-3,4

OpenStudy (anonymous):

Help give medal

pooja195 (pooja195):

\[\huge~\rm~x-1=\frac{ -4 }{ 3 }x-8\] Solve for x I plugged equation 1 into equation 2

OpenStudy (anonymous):

So then if you substitute you get -16/3 but it not equal to -8

pooja195 (pooja195):

Now quite... \[\huge~\rm~x-1+4/3=\frac{ -4 }{ 3 }x+4/3-8\] \[\huge~\rm~\frac{ 7 }{ 3 }x-1=-8\] \[\huge~\rm~\frac{ 7 }{ 3 }x-1+1=-8+1\] \[\huge~\rm~\frac{ 7 }{ 3 }x=-7\] Now solve for x

OpenStudy (taylor0402):

i think you have all the help you need she seems to be doing great

OpenStudy (taylor0402):

achuley perfect

pooja195 (pooja195):

Thanks! =)

OpenStudy (anonymous):

If you plug 4 for x right because u use the first problem

pooja195 (pooja195):

Now we are only solving for x

pooja195 (pooja195):

For right now

OpenStudy (taylor0402):

yea you are very good at math

OpenStudy (anonymous):

But then how do use substition method

pooja195 (pooja195):

\[\huge~\rm~\frac{ 7 }{ 3 }x \div \frac{ 7 }{ 3 }x=-7 \div \frac{ 7 }{ 3 }x\]

pooja195 (pooja195):

Did you look at the steps I wrote out above?

OpenStudy (anonymous):

Don't you plug the first plug 1 into 2

pooja195 (pooja195):

It doesnt matter how you plug them in....you sill get the same answer

pooja195 (pooja195):

\[\huge~\rm~x=-7 \div \frac{ 7 }{ 3 }x\]

OpenStudy (anonymous):

Yea you added-4/3 to both sides

pooja195 (pooja195):

We are way past that...

OpenStudy (anonymous):

I got -1/3 but

OpenStudy (anonymous):

For x

pooja195 (pooja195):

Thats incorrect What is -7 divided by 7/3 ?

OpenStudy (anonymous):

Oh it's -3

OpenStudy (anonymous):

Oh ok I get it know so then it is d right

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

My bad answer d was suppose to be -3,-4

OpenStudy (anonymous):

Does that change anything

pooja195 (pooja195):

D is correct correct =)

pooja195 (pooja195):

What do you mean by "Does it change anything?"

OpenStudy (anonymous):

But the 4 in d I forgo to put - sign in front of the 4

OpenStudy (anonymous):

That's why I was asking if it changes anything

pooja195 (pooja195):

We know that x=-3 right We can substiitue the value of x into y=x-1 \[\huge~\rm~ y=(-3)-1\] \[\huge~\rm~ -4=(-3)-1\] So -3-1=-4 therefore y=-4

OpenStudy (anonymous):

@taylor0402 thank you anways

OpenStudy (anonymous):

Oh ok thank so much you been do patient with me @pooja195

OpenStudy (anonymous):

It d then

pooja195 (pooja195):

haha no problem ^-^ Our goal is to help users understand =) their question patience is the virtue! Yes correct it is D :)